OpenStudy (anonymous):

solve for x square root 5x+3=8

OpenStudy (anonymous):

\[\sqrt{5x+3}=8\] \[(\sqrt{5x+3})^{2}=(8)^{2}\]

OpenStudy (anonymous):

5x=61 x=61/5

OpenStudy (anonymous):

There you go :)